OpenStudy (anonymous):

Alright, so i need help with finding out what i did wrong on this...? Teacher said i need more but dont KNOW WHAT TO ADD! She said add examples and a few changes. One teaching example for each of the four questions, which includes step by step instructions written in complete sentences. 1.If I know the lengths of two sides of a right triangle, how do I find the third? 2.Could I find the two missing side lengths of a right triangle if I only know one side length and one angle measure (other than the 90 degree angle)? 3.Could I find the two missing angle measures if I know some of the sid

OpenStudy (anonymous):

1. Alright so to start off, in order to find the last side of the triangle, we will be using the Pythagorean Theorem, which is an equation used to find the third sides of triangles, most of the time, a hypotenuse. The equation it uses is a2+b2=c2. So basically say you have the two smaller sides of the triangle with there measurements. Then you take and plug in there numbers into the equation so for instance one side is 2 and the other side is 3. That would be 2squared+3squared=csquared. C is the missing segment. So in order to get C it would be 2*2+3*3=10squared which would equal 3.2. Another way to use this to your advantage to get one of the smaller sides is if is was c2-b2=a2 which basically you would do what you did on the first one except backwards. Just plug in the numbers and everything except this time your going to subtract one of the sides from the hypotenuse to get your other smaller side. 2. Yes using sine, cosine, and tangent. Yes, using the Law of Sines you can. The way you use the Law Of Sines is using the the three properties, Sin, Cosine, and Tangent if its a 30-60-90 triangle. Label the 30 degree angle and then label the side opposite as 1/2 to show you can find the Hypotenuse: h/sin(90)=0.5/sin(30) and then show the second side which would be x/sin(60)=0.5/sin(30). 3. Yea because as long as you know at least two sides then you can find missing angles. Yes you can because as long as at least two sides are known, then you can apply them to the Law of Sines. Then form there you can use the Pythagorean Theorem to find the Hypotenuse. 4. A special right triangle is where you have two angles, 45 and 45, then you have 90. It is special because of the 3:4:5 ratio and because it helps allow quick calculations of lengths in geometry problems.

OpenStudy (anonymous):

Please help as its holding me back >.>

OpenStudy (anonymous):

..1. The easiest way would be with the Pythagorean theorem. A^2+B^2=C^2. So if you know any two sides, you can find the third side pretty easily. (It helps to have a calculator to do the squares and square roots). 2. Yes. The Law of Sines has the ratios for all the sides given the angles, or vice versa. There are three sides and three angles (of which one angle is always 90 degrees--that's what makes it a right triangle, it has a 'right angle' in it). So it comes down to three lengths of sides and two angles. If you know any two of these sides or angles, you can calculate the others using the law of sines. 3. Yes. See answer no. 2 4. I don't know what a 'special' right triangle is. What makes them special?

OpenStudy (anonymous):

@TrashMan914

OpenStudy (anonymous):

Looks like seiga here gave you an answer. Does that not help?

OpenStudy (phi):

your question was cut off, so only part of 3 and none of 4 is given. but I would add the law of cosines for question 3. the special triangles are 30-60-90 and 45-45-90 your write up on question 4 is It is special because of the 3:4:5 ratio and because it helps allow quick calculations of lengths in geometry problems. does not make sense. they are special because you can write the lengths of the sides by inspection, and do not need a calculator.